the price elasticity of rice is estimated to be -0.4 and the income elasticity is 0.8. At a price of $0.40 per pound and a per capital income of $20,000, the demand for rice is 50million tons per year.

a. is rice an inferior good, a necessity or a luxury? Explain

b. if per capital income increases to $20,500, what will be the quantity demanded of rice?


1
Expert's answer
2021-02-05T13:33:33-0500

The quantity of rice demanded is =  If per capita income increases to $20,500, the change in quantity demanded of rice will be 0.8"\times" "\frac{(20,500 - 20,000)}{20000}" = 0.02 or +2%.

If the price of rice increases to $0.41 per pound and income per capita remains at $20,000, the change in quantity demanded is -0.4"\times" "\frac{(0.41 - 0.40)}{0.40} = -1" , so the new quantity demanded is Qd = 50 - 1 = 49 million tons.